OpenStudy (anonymous):

Simplify: (a^4/3)^(-2)

OpenStudy (anonymous):

a^(4/3 * -2) 4/3 * -2 is -8/3 a^(-8/3) is the answer

OpenStudy (lalaly):

nooo \[(\frac{a^4}{3})^{-2}=(\frac{3}{a^4})^2 = \frac{9}{a^8}\]

OpenStudy (akshay_budhkar):

when there is a negative radical then the number goes in the denominator like in this case we have (3/a^4)=9/a^8

OpenStudy (akshay_budhkar):

oh lana did it before :D

OpenStudy (lalaly):

:D rita is givin wrong answers evrywhere:S

OpenStudy (akshay_budhkar):

yea she has....

OpenStudy (anonymous):

Sorry I thought a to the power of 4/3

OpenStudy (anonymous):

Thats why there is so many off us to help out, though I am not much help

OpenStudy (akshay_budhkar):

no problem rita!! mistakes do happen!! and ehemze dont worry you will also help in sometime